# SafeGen: Mitigating Unsafe Content Generation in Text-to-Image Models

### v1

@@ 3 Analysis on Adversarial Prompts
In this section, we briefly introduce existing countermeasures against NSFW image generation, including external [11, 12] and internal [13, 14] measures. Then, we reveal insufficient protection provided by existing defense methods under adversarial prompts, which motivates us to design a new text-agnostic defense framework.
3.1 Defenses Against NSFW Image Generation
@@
Existing defenses against pornographic image generation mainly focus on employing external defenses to filter harmful content and internal defenses to suppress sexually explicit concepts. External text- and image-based safety filters [10, 11] are widely adopted by commercial service providers [4, 5] and open-source model platforms, e.g., HuggingFace [38]. These plug-in filters either deny the textual input containing pornographic words [10], or obstruct the resulting image into black upon detecting pornographic output [11], as depicted in Figure 5. Hence, T2I models may be enhanced to be resistant to the influence of unsafe prompts. External detection methods also include Stable Diffusion 2.1 (SD-V2.1) [12], since it is retrained on cleansed data, where NSFW information is censored by external safety filters. The internal defenses encompass safe latent diffusion (SLD) [13] and erased stable diffusion (ESD) [14], which are all text-dependent. SLD [13] prohibits a bag of negative concepts (e.g., naked body) and enhances the classifier-free guidance with a new conditioned diffusion item to shift away from unsafe regions. ESD [14] modifies the SD model to suppress sexual parts of input text (e.g., “a nude man” to “a man”). However, a noteworthy research question arises: Are existing protections enough in preventing unsafe image generation?
@@
Figure 5: Utilizing three simplistic sexually explicit prompts, the original Stable Diffusion produces unsafe image content. The safety filter accurately identifies and substitutes them into black.
3.2 Impact of Adversarial Prompts
@@
Unfortunately, our analysis unveils a worrisome picture. Adversarial prompts [16, 17] are shown to drive T2I models to generate NSFW content under existing defenses, as shown in Figure 6. Safety filters fail to filter inappropriate text and prevent NSFW image generation. SD-V2.1, though being retrained on filtered data, still generates NSFW images. The root cause is that inherent under-generalization of detection models [11, 39] leads to undetected errors after images created and unfiltered pornographic samples in the censored training dataset. ESD [14] neutralizes sexual concepts such as “nudity” to “[blank]” by fine-tuning the parameters of cross-attention layers of Stable Diffusion. In this way, unseen sexual concepts with embedding-level proximity to known sexual concepts (e.g., “naked, porn, sexy”) may also be suppressed thanks to the well-trained CLIP text encoder [29]. However, it is shown that ESD is still vulnerable to adversarial prompts. The reason lies in concepts that seem to be innocent but connote sexual meanings. Taking the prompt (a) from the I2P dataset [18] as an instance, the names of porn stars, “M** D**” and “C** M**”, are dissimilar to those suppressed explicit words at the embedding level, inducing NSFW image generation. Due to a similar reason that adversarial prompts differ from the predefined unsafe concepts at the embedding level, SLD [13] is also enticed by adversarial prompts to generate erotic images. Based on the above findings, we summarize existing defenses as follows:
$\bullet$ External defenses. Text- and image-based safety filters [10, 11] are efficient to deploy for filtering NSFW content [12]. However, detection models face the inherent challenge of under-generalization, exhibiting inadequacies in filtering adversarial prompts and explicit images. $\bullet$ Internal defenses. SLD [13] and ESD [14] aim to enhance the endogenous compliance of T2I models. Nevertheless, these text-dependent methods encounter difficulties in seemingly unrelated but sexually connoted words, which are difficult to enumerate and can induce visually explicit representations within the T2I models.
@@
Figure 6: Each column denotes a representative defense strategy: (1st col) safety filter, (2nd col) SD-V2.1, (3rd col) SLD, and (4th col) ESD. From prompt (a) to (c), each row corresponds to an adversarial prompt (listed in Appendix A), which can compromise all these latest defense strategies and allure Stable Diffusion to generate unsafe images.
The fact that existing defense countermeasures fail to resist adversarial prompts inspires us to design SafeGen, a text-agnostic protection framework that is robust to adversarial prompts.
4 Design of Text-Agnostic SafeGen
4.1 Overview
@@
Figure 7: The impact of overall quality and semantics of generated images wi/wo modifying the unconditionally vision-only diffusion process. The original Stable Diffusion (1st row); Stable Diffusion with the vision-only process modified (2nd row).
@@
Figure 8: Diagram of governing the self-attention layers. The data preparation includes <nude,mosaic,benign> image triplets (in the blue box), where benign $x^{b}$ and nude $x^{n}$ images as input, along with the mosaic output $x^{m}$. The adjustment process for self-attention layers involves iteratively injecting random noise into the latent space of each image, followed by the denoising U-Net purifying the noisy latent $T$ times. Consequently, the visually explicit latent representations are obscured as $z^{m}_{0}$, while the matrices $\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}$ of self-attention layers preserve the ability to represent benign visual latent $z^{b}_{0}$.
4.2 Rationale Behind Text-Agnostic Design
@@
In revisiting the generation mechanism of T2I models, i.e., classifier-free guidance mentioned in §2.2, we verify that managing its unconditionally vision-only denoising diffusion process $\epsilon_{\mathtt{U}}(z_{t},t)$ alone can significantly impact the overall quality and semantics of the resulting images. Specifically, as shown in Figure 7, we perform a comparative analysis to examine the impact of modifying the unconditional process within the classifier-free guidance. While the conditional guidance term $\epsilon_{\mathtt{U}}(z_{t},c,t)$ keeps an identical text embedding $c$, the images generated by the modified SD model (the 2nd row) are distinct from the original set (the 1st row). The semantics of images in the 2nd row are hard to interpret and drastically deviate from the user’s desired output in the 1st row. A diversity of images is generated despite identical textual prompts due to initial sampling variations in the latent distribution with disparate random seeds. The goal of the unconditional process is to make the generated images resemble real image distributions, which is achieved by iteratively purifying the noisy latent into cleaner latent. However, if we modify the denoising U-Net so that it is unable to clear up visually explicit latent representations, then the guidance provided by the NSFW text conditions becomes ineffective. Hence, a crucial inquiry is how to autonomously obscure or corrupt nude areas during the denoising diffusion process, which serves as a foundation for ensuring the safety of any generated image in a text-agnostic manner.
4.3 Governing Vision-Only Self-Attention Layers
@@
We aim to enable the unconditionally text-agnostic denoising diffusion process to autonomously corrupt sexually explicit regions. Considering the convolutional and self-attention layers involved in this process, we choose the self-attention mechanism due to its multifaceted advantages over CNNs as outlined in §2.3.2. In particular, its proficiency in comprehending the association among pixels and their overall semantics is useful for locating explicit regions. Our empirical experiments also justify that solely modifying self-attention layers would outperform optimizing all text-independent modules for this objective, with the same hyperparameters given in Appendix C.
@@
Figure 8 presents our scheme to regulate the $\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}$ matrices of SD model’s self-attention layers from original $\mathbf{W}$ to protected $\mathbf{W^{*}}$, using <nude, censored, benign> image triplets. The data preparation employs a mosaic neural network [40] to automatically mask a batch of pornographic images $x^{n}$ with thick mosaic to derive the mosaic images $x^{m}$. As our model editing involves corrupting human nudity representations, which may impact the ability of benign human-oriented image generation, we randomly sample everyday benign photos $x^{b}$ from Human Detection Dataset [41] as benign counterparts. In effect, with merely 100 <nude, censored, benign> image triplets, the self-attention layers can swiftly unlearn pornographic representations and effectively corrupt the latent’s explicit regions.
@@
Before adjusting self-attention layers, SD model’s encoder $\mathcal{E}$ transforms the <nude, censored, benign> triplets <$x^{n},x^{m},x^{b}$> into clean latent representations <$z^{n}_{0},z^{m}_{0},z^{b}_{0}$>. Then the DDPM noise scheduler [1] iteratively injects noise $\epsilon^{n}_{t}$ and $\epsilon^{b}_{t}$ into the images at each time step $t$, forming <$z^{n}_{t},z^{m}_{t},z^{b}_{t}$> and resulting in the final noisy <$z^{n}_{T},z^{m}_{T},z^{b}_{T}$> triplets. It is noteworthy that we let the DDPM scheduler inject the same noise $\epsilon^{n}_{t}$ on the nude and mosaic latent, which is related to the loss function Equation (4) (detailed in Appendix B). Subsequently, in the denoising diffusion process, we always inject the cross-attention layers (as outlined in §2.3.1) with a piece of blank textual information "". This ensures that self-attention layers can unconditionally remove pornographic latent representations from its attentive matrices $\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}$ step by step via optimization, cutting off the associations between sexually-related text and nudity vision. The blank injection operation also renders Equation (2) to $\epsilon_{\mathtt{U^{*}}}(z_{t},t)$ as employed in Equation (4), (5). After $T$ timesteps, the U-Net is expected to gradually purify visually-nude noisy latent to censored latent $z^{n}_{T}\rightarrow z^{m}_{0}$, while ensuring visually-benign latent is restored to its originally clean latent $z^{b}_{T}\rightarrow z^{b}_{0}$. To realize this objective, our two loss function terms $\mathcal{L}_{m}$ (Loss mosaic) and $\mathcal{L}_{p}$ (Loss preservation) are expressed as follows:
$\mathcal{L}_{m}=\sum_{t=0}^{T}{\epsilon_{\mathtt{U^{*}}}(z^{n}_{t},t)}-(\hat{z}^{n}_{T}-\hat{z}^{m}_{T}+\sum_{t=0}^{T}{\epsilon^{n}_{t}})$ (4)
$\mathcal{L}_{p}=\sum_{t=0}^{T}{\epsilon_{\mathtt{U^{*}}}(z^{b}_{t},t)}-\sum_{t=0}^{T}{\epsilon^{b}_{t}}$ (5)
@@
where minimizing $\mathcal{L}_{m}$ encourages self-attention layers to remove nude representations, i.e., projecting them to latent covered with thick mosaic. We detail the proof of Equation (4) in Appendix B. Scaling down $\mathcal{L}_{p}$ forces these layers to maintain benign image representation quality and avoid parameter shifts. More specifically, $\epsilon^{k}_{t}\sim\mathcal{N}(0,I^{2}),k\in[n,b,m]$. Each $\epsilon^{k}_{t}$ added on the original latent $z^{k}_{0}$ is predefined by the DDPM scheduler. In other words, $\sum_{t=0}^{T}\epsilon^{k}_{t}$ denotes their summation for the entire noise injection process, and $z^{k}_{T}=z^{k}_{0}+\sum_{t=0}^{T}\epsilon^{k}_{t}$. Similarly, $\sum_{t=0}^{T}\epsilon_{\mathtt{U^{*}}}(z^{k}_{t},t)$ denotes the aggregate noise predicted by the U-Net $\mathtt{U^{*}}$ with adjusted self-attention layers. Ideally, this term equals $\sum_{t=0}^{T}\epsilon^{k}_{t}$.
$\underset{\mathbf{W^{*}}}{\text{min}}(\lambda_{m}\mathcal{L}_{m}+\lambda_{p}\mathcal{L}_{p})$ (6)
@@
The two objectives in Equation (6) can be optimized jointly via AdamW optimizer [42]. Our experiments demonstrate the settings of $\lambda_{m}:$ 0.1, $\lambda_{p}:$ 0.9 can realize the ideal performance of both nudity removal and benign preservation, as shown in Figure 9. Additionally, we provide a more detailed comparison between SafeGen and existing methods in terms of mitigating sexually explicit generation (see Appendix D, Figure 11) while preserving the ability to generate high-fidelity images of various non-explicit categories (see Appendix F, Figure 13).
4.4 System Integration

### v2

@@ 3. Related Work & Motivation
In this section, we review existing attacks that induce T2I models to produce unsafe content, along with countermeasures to defend explicit generation. These defenses include external (huggingface_safety_checker, ; SD-2-1, ) and internal (schramowski2023safe, ; gandikota2023erasing, ) measures. Then, we reveal insufficient protection provided by existing defense methods under adversarial prompts, which motivates us to design a new text-agnostic defense framework.
3.1. Attacks on Text-to-Image (T2I) Models
The susceptibility of T2I models to generating NSFW content, particularly sexual explicitness, has been a significant concern (AI_porn, ; brack2023mitigating, ; bird2023typology, ). This issue has spurred investigations into various attack vectors targeting these models, such as red-teaming the SD model for unsafe image generation (rando2022red, ) through reverse engineering its safety filter mechanism. Moreover, adversarial prompts (tsai2023ring, ; yang2023sneakyprompt, ; gao2023evaluating, ) have been crafted to manipulate T2I models into producing unsafe images while evading detection. For instance, Ring-A-Bell (tsai2023ring, ) tailors adversarial textual inputs that are conceptually close to the target yet contain nonsense words. Gao et al. (gao2023evaluating, ) introduces a word-level similarity constraint to mimic realistic human errors, e.g., typo, glyph, and phonetic mistakes. SneakyPrompt (yang2023sneakyprompt, ) employs a reinforcement learning-based search approach to create adversarial prompts that preserve NSFW semantics, effectively bypassing safety mechanisms in SD models. Another vulnerability is the reliance of T2I models on large datasets, which may be susceptible to poisoning attacks. Adversaries can release poisoned text-image data online (wu2023proactive, ), which is then inadvertently collected by data trainers, leading to potential unethical outputs from T2I models.
3.2. Defenses Against Explicit Generation
@@
The generation of sexually explicit content has highlighted the critical need to regulate T2I models. Current strategies focus on employing external defenses to filter harmful content and internal defenses to suppress sexually explicit concepts. External text- and image-based safety filters (text_safety_classifier, ; huggingface_safety_checker, ) are widely adopted by commercial service providers (midjourney, ; dalle2_openai, ) and open-source model platforms, e.g., HuggingFace (huggingface, ). These plug-in filters either deny the textual input containing explicit words (text_safety_classifier, ), or obstruct the resulting image into black upon detecting sexually explicit output (huggingface_safety_checker, ), as depicted in Figure 5. Hence, T2I models may be enhanced to be resistant to the influence of unsafe sexual prompts. External detection methods also include Stable Diffusion 2.1 (SD-V2.1) (SD-2-1, ), since it is retrained on cleansed data, where NSFW information is censored by external safety filters. The internal defenses encompass safe latent diffusion (SLD) (schramowski2023safe, ) and erased stable diffusion (ESD) (gandikota2023erasing, ), which are all text-dependent. SLD (schramowski2023safe, ) prohibits a bag of negative concepts (e.g., naked body) and enhances the classifier-free guidance with a new conditioned diffusion item to shift away from unsafe regions. ESD (gandikota2023erasing, ) modifies the SD model to suppress sexual parts of input text (e.g., “a nude man” to “a man”). However, a noteworthy research question arises: Are existing protections enough in preventing unsafe image generation?
@@
Figure 5. Utilizing three simplistic sexually explicit prompts, the original Stable Diffusion produces unsafe image content. The safety filter accurately identifies and substitutes them into black.
3.3. Impact of Adversarial Prompts
@@
Unfortunately, our analysis unveils a worrisome picture. Adversarial prompts (yang2023sneakyprompt, ; qu2023unsafe, ) are shown to drive T2I models to generate sexually explicit content under existing defenses, as shown in Figure 6. Safety filters fail to filter inappropriate text and prevent unsafe image generation. SD-V2.1, though being retrained on filtered data, still generates NSFW images. The root cause is that inherent under-generalization of detection models (huggingface_safety_checker, ; LAION_filter, ) leads to undetected errors after images created and unfiltered pornographic samples in the censored training dataset. ESD (gandikota2023erasing, ) neutralizes sexual concepts such as “nudity” to “[blank]” by fine-tuning the parameters of cross-attention layers of Stable Diffusion. In this way, unseen sexual concepts with embedding-level proximity to known sexual concepts (e.g., “naked, porn, sexy”) may also be suppressed thanks to the well-trained CLIP text encoder (CLIP, ). However, it is shown that ESD is still vulnerable to adversarial prompts. The reason lies in concepts that seem to be innocent but connote sexual meanings. Taking the prompt (a) from the I2P dataset (I2P, ) as an instance, the names of porn stars, “M** D**” and “C** M**”, are dissimilar to those suppressed explicit words at the embedding level, inducing sexually explicit image generation. Due to a similar reason that adversarial prompts differ from the predefined unsafe concepts at the embedding level, SLD (schramowski2023safe, ) is also enticed by adversarial prompts to generate erotic images. Based on the above findings, we summarize existing defenses as follows:
@@
Figure 6. Each column denotes a representative defense strategy: (1st col) safety filter, (2nd col) SD-V2.1, (3rd col) SLD, and (4th col) ESD. From prompt (a) to (c), each row corresponds to an adversarial prompt (listed in Appendix A), which can compromise all these latest defense strategies and allure Stable Diffusion to generate unsafe images.
Our Approach. Unlike prior countermeasures, SafeGen makes the first attempt to remove representations of visually sexual content from Stable Diffusion in a text-agnostic manner. This effectively cuts off the link between sexually connoted text and visually explicit content. In addition, SafeGen retains the capability for benign image generation and can seamlessly integrate with existing defense techniques.
4. Design of Text-Agnostic SafeGen
4.1. Overview
@@
Figure 7. The impact of overall quality and semantics of generated images wi/wo modifying the unconditionally vision-only diffusion process. The original Stable Diffusion (1st row); Stable Diffusion with the vision-only process modified (2nd row).
@@
Figure 8. Diagram of governing the vision-only self-attention layers. The data preparation includes <nude,mosaic,benign> image triplets (in the blue box), where benign $x^{b}$ and nude $x^{n}$ images as input, along with the mosaic output $x^{m}$. The adjustment process for self-attention layers involves iteratively injecting random noise into the latent space of each image, followed by the denoising U-Net purifying the noisy latent $T$ times. Consequently, the visually explicit latent representations are obscured as $z^{m}_{0}$, while the matrices $\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}$ of self-attention layers preserve the ability to represent benign visual latent $z^{b}_{0}$.
4.2. Rationale Behind Text-Agnostic Design
@@
In revisiting the generation mechanism of T2I models, i.e., classifier-free guidance mentioned in §2.2, we verify that managing its unconditionally vision-only denoising diffusion process $\epsilon_{\mathtt{U}}(z_{t},t)$ alone can significantly impact the overall quality and semantics of the resulting images. Specifically, as shown in Figure 7, we perform a comparative analysis to examine the impact of modifying the unconditional process within the classifier-free guidance. While the conditional guidance term $\epsilon_{\mathtt{U}}(z_{t},c,t)$ keeps an identical text embedding $c$, the images generated by the modified SD model (the 2nd row) are distinct from the original set (the 1st row). The semantics of images in the 2nd row are hard to interpret and drastically deviate from the user’s desired output in the 1st row. A diversity of images is generated despite identical textual prompts due to initial sampling variations in the latent distribution with disparate random seeds. The goal of the unconditional process is to make the generated images resemble real image distributions, which is achieved by iteratively purifying the noisy latent into cleaner latent. However, if we modify the denoising U-Net so that it is unable to clear up visually explicit latent representations, then the guidance provided by the unsafe text conditions becomes ineffective. Hence, a crucial inquiry is how to autonomously obscure or corrupt nude areas during the denoising diffusion process, which serves as a foundation for ensuring the safety of any generated image in a text-agnostic manner.
4.3. Governing Self-Attention Layers
@@
We aim to enable the unconditionally text-agnostic denoising diffusion process to autonomously corrupt sexually explicit regions. Considering the convolutional and self-attention layers involved in this process, we choose the self-attention mechanism due to its multifaceted advantages over CNNs as outlined in §2.3.2. In particular, its proficiency in comprehending the association among pixels and their overall semantics is useful for locating explicit regions. Our empirical experiments also justify that solely modifying self-attention layers would outperform optimizing all text-independent modules for this objective, with the same hyperparameters given in Appendix B.
@@
Figure 8 presents our scheme to regulate the $\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}$ matrices of SD model’s self-attention layers from original $\mathbf{W}$ to protected $\mathbf{W^{*}}$, using <nude, censored, benign> image triplets. The data preparation employs a mosaic neural network (anti_deepnude, ) to automatically mask a batch of pornographic images $x^{n}$, which are from the NSFW dataset (large_NSFW_image, ), with thick mosaic to derive the mosaic images $x^{m}$. As our model editing involves corrupting human nudity representations, which may impact the ability of benign human-oriented image generation, we randomly sample everyday benign photos $x^{b}$ from Human Detection Dataset (human_dataset, ) as benign counterparts. In effect, with merely 100 randomly selected <nude, censored, benign> image triplets, the self-attention layers can swiftly unlearn pornographic representations and effectively corrupt the latent’s explicit regions.
@@
Before adjusting self-attention layers, SD model’s encoder $\mathcal{E}$ transforms the <nude, censored, benign> triplets <$x^{n},x^{m},x^{b}$> into clean latent representations <$z^{n}_{0},z^{m}_{0},z^{b}_{0}$>. Then the DDPM noise scheduler (ho2020denoising, ) iteratively injects noise $\epsilon^{n}_{t}$ and $\epsilon^{b}_{t}$ into the images at each time step $t$, forming <$z^{n}_{t},z^{m}_{t},z^{b}_{t}$> and resulting in the final noisy <$z^{n}_{T},z^{m}_{T},z^{b}_{T}$> triplets. It is noteworthy that we let the DDPM scheduler inject the same noise $\epsilon^{n}_{t}$ on the nude and mosaic latent, which is related to the loss function Equation (4) (detailed in Appendix C). Subsequently, in the denoising diffusion process, we always inject the cross-attention layers (as outlined in §2.3.1) with a piece of blank textual information "". This ensures that self-attention layers can unconditionally remove pornographic latent representations from its attentive matrices $\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}$ step by step via optimization, cutting off the associations between sexually-related text and nudity vision. The blank injection operation also renders Equation (2) to $\epsilon_{\mathtt{U^{*}}}(z_{t},t)$ as employed in Equation (4), (5). After $T$ timesteps, the U-Net is expected to gradually purify visually-nude noisy latent to censored latent $z^{n}_{T}\rightarrow z^{m}_{0}$, while ensuring visually-benign latent is restored to its originally clean latent $z^{b}_{T}\rightarrow z^{b}_{0}$. To realize this objective, our two loss function terms $\mathcal{L}_{m}$ (Loss mosaic) and $\mathcal{L}_{p}$ (Loss preservation) are expressed as follows:
(4) $\mathcal{L}_{m}=\sum_{t=0}^{T}\left\|\epsilon_{\mathtt{U^{*}}}(z^{n}_{t},t)-(\hat{z}^{n}_{T}-\hat{z}^{m}_{T}+\sum_{t=0}^{T}{\epsilon^{n}_{t}})\right\|_{2}^{2}$
(5) $\mathcal{L}_{p}=\sum_{t=0}^{T}\left\|\epsilon_{\mathtt{U^{*}}}(z^{b}_{t},t)-\epsilon^{b}_{t}\right\|_{2}^{2}$
@@
where minimizing $\mathcal{L}_{m}$ encourages self-attention layers to remove nude representations, i.e., projecting them to latent covered with thick mosaic. We detail the proof of Equation (4) in Appendix C. Scaling down $\mathcal{L}_{p}$ forces these layers to maintain benign image representation quality and avoid parameter shifts. More specifically, $\epsilon^{k}_{t}\sim\mathcal{N}(0,I^{2}),k\in[n,b,m]$. Each $\epsilon^{k}_{t}$ added on the original latent $z^{k}_{0}$ is predefined by the DDPM scheduler. In other words, $\sum_{t=0}^{T}\epsilon^{k}_{t}$ denotes their summation for the entire noise injection process, and $z^{k}_{T}=z^{k}_{0}+\sum_{t=0}^{T}\epsilon^{k}_{t}$. Similarly, $\sum_{t=0}^{T}\epsilon_{\mathtt{U^{*}}}(z^{k}_{t},t)$ denotes the aggregate noise predicted by the U-Net $\mathtt{U^{*}}$ with adjusted self-attention layers. Ideally, this term equals $\sum_{t=0}^{T}\epsilon^{k}_{t}$.
(6) $\underset{\mathbf{W^{*}}}{\text{min}}(\lambda_{m}\mathcal{L}_{m}+\lambda_{p}\mathcal{L}_{p})$
@@
The two objectives in Equation (6) can be optimized jointly via AdamW optimizer (loshchilov2017decoupled, ). Our experiments demonstrate the settings of $\lambda_{m}:$ 0.1, $\lambda_{p}:$ 0.9 can realize the ideal performance of both nudity removal and benign preservation, as shown in Figure 9. Additionally, we provide a more detailed comparison between SafeGen and existing methods in terms of mitigating sexually explicit generation (see Appendix D, Figure 15) while preserving the ability to generate high-fidelity images of various non-explicit categories (see Appendix F, Figure 17).
4.4. System Integration